# ASE: Practical Acoustic Speed Estimation Beyond Doppler via Sound Diffusion Field

### v1

@@ 4.2 Transmission Scheme
In this part, we will elaborate our transmission scheme design. We will then describe the channel estimation principles.
@@
Single-Speaker OTDM: We stack two orthogonal signals as a complex signal and pass it into the I/Q modulator for transmission. Specifically, we can treat the orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$ as the real and imaginary part of a complex signal, i.e., $s(t)=s_{1}(t)+js_{2}(t)$. To delay $s_{2}(t)$, we shift the sequence by padding zeros to the front of $s_{2}(t)$. Practically, however, a speaker can only transmit a single sequence of real signals. To transmit the resultant complex signal over a speaker, we pass it through an I/Q modulator to convert it into a real one. As illustrated in Fig. 6, we additionally multiply $s_{1}(t)$ and $s_{2}(t)$ with a carrier signal of orthogonal phase and obtain the modulated real signal: $s(t)=s_{1}(t)\cos(2\pi f_{c}t)-s_{2}(t)\sin(2\pi f_{c}t)$. With such modulation, the two components remain orthogonal to each other, while the modulated signal can be transmitted through a single speaker. At Rx, the received signal is first demodulated by the carrier signal with a low-pass filter (LPF). We can then separate the overlapped signal and estimate the CSI for each sequence at the receiver’s side, as shown in Fig. 6. By doing so, we can obtain two separated signals $r_{1}(t)$ and $r_{2}(t)$. Then channel estimation will be performed as follows.
@@
Figure 8: Peak prominence weighting examples. Top: Violin plots of the distribution of the delays of the first peak. Middle: Original weight distribution, where the circle size indicates the weights. The red dashed line represents the fitted $\sigma(\tau)$ Bottom: Adjusted prominence weights.
@@
Channel Estimation: To capture the channel properties, we generate the Kasami sequence $s(t)$ of length $N_{s}$ as the probe signal and transmit it over a speaker. The signal interacts with the environment before arriving at the microphone, where it undergoes scatters in the sound diffusion field. At the receivers end, we acquire $r(t)$ and separate them into $r_{1}(t)$ and $r_{2}(t)$, respectively. The CIR $h(t)$ is then estimated as the correlation between them, i.e., $h_{i}(t)=r_{i}(t)\ast s(t),i\in\{1,2\}$. Then CSI $H_{i}(f,t)$ is obtained by transforming $h_{i}(t)$ into the frequency domain via FFT. $H_{1}(f,t)$ and $H_{2}(f,t)$ will be then merged using the OTDM scheme to acquire the boosted channel estimation $H(f,t)$, as illustrated in §2.
4.3 Speed Estimation
@@
Weighted Subcarrier Combining: Proposed model accounts for all multipath components for speed estimation. Recall Eq. (13), however, the theoretical model estimates speed from one single subcarrier. Given multiple subcarriers available, we can further embrace frequency diversity to boost speed estimation. Instead of performing speed estimation on each subcarrier and fusing the estimates, we propose to properly combine $\psi_{G}(f,t)$ calculated on each subcarrier to enhance the SNR of the speed signal. Formally, we aim to obtain the boosted ACF as $\hat{\psi}_{G}(\tau)=\sum w(f)\psi_{G}(f,\tau)$, where
$\begin{gathered}\max_{w}\quad\hat{\psi}_{G}(\tau_{s})=\sum_{i}w(f_{i})\psi_{G}(f_{i},\tau_{s_{i}}),\\ \begin{aligned} &\text{s.t.}\quad\sum_{i}w(f)=1,w(f)\geq 0,\forall f\in\mathbb{F},\\ \end{aligned}\end{gathered}$ (15)
@@
Frequency Alignment: The weighted combining will enhance the speed signals and reduce the noise, only if the signals are synchronized. However, significant subcarrier frequency variations misalign the ACFs across subcarriers, particularly the first peak, for the same speed. Take two subcarriers on 20 kHz and 24 kHz as an example. Considering a speed $v=0.5m/s$, the peaks will appear at a delay $\tau=x_{0}\lambda(f)/2\pi v$. The wavelength $\lambda(f)$ depends on the carrier frequency, and will be 1.715 cm for 20 kHz and 1.429 cm for 24 kHz, leading to two considerably different delays for the first peaks in the respective ACFs. Therefore, we need to eliminate the subcarrier frequency differences and align all the ACFs. We scale the ACF in the lag domain and align them with respect to the first peaks. Particularly, we choose a virtual frequency as a reference, denoted as $f_{ref}$, and scale the ACF on all subcarriers to the same $f_{ref}$, i.e., $\psi(f,\tau)\to\psi(f_{ref},\tau^{\prime})$. Recall Eq. (14), we obtain the the aligned first peak as, $\tau_{s}^{\prime}={f}/{f_{ref}}\cdot\tau_{s}$. Then the weighted combining can be done on the aligned ACF.
Figure 9: Experiment Setting
5 Evaluations
5.1 Implementation
@@
Hardware: We implement ASE with programmable smart speaker prototype, i.e., a microphone [5], and speaker[2]. Notably, we only use one microphone channel in our experiments. The microphone and the speaker are co-located.
@@
Software: We implement the algorithms of ASE in Python and Matlab. Specifically, we develop a sound playing and recording program with Python and implement the pipeline of processing algorithms with Matlab. We apply 1-s window to compute ACF with a step of 0.1s. We perform motion detection, based ZCC, to determine whether a user is moving and perform speed estimation when motion is detected. Before weighted subcarrier combining, we also perform outlier detection to sift out some abnormal ACFs, which occasionally appears as either a near-linear trend or a zig-zag spikes.
@@
As shown in Fig. 9, we mainly conduct our experiment in a 4 $\times$ 4m room equipped with the OptiTrack [4], which serves as the ground truth. We derive speed from the solved whole-body skeleton coordinates from OptiTrack. We have gained IRB approval by our university for data collection. In total, we have collected over 700 minutes of moving data traces. We evaluate the performance of ASE under various settings. We use Mean Speed Error (MSE) and detection rate (i.e., how often ASE can reliably detect a speed) as the main evaluation metrics.
Figure 10: MSE-DW
Figure 11: MSE-CW
Figure 12: MSE-RW
Figure 13: Detection Rate vs. Walk Means
Figure 14: CDF vs. Distances
Figure 15: CDF vs. Walk Means
Figure 16: MSE of NLOS Case
5.2 Overall Performance
@@
Direct Walk (DW): We place the device at the center of one wall in the room. Participants are asked to walk towards the audio device from 1m to 3.5m. We calculate the mean speed error in Fig. 16 and detection rate in Fig. 16. Our results demonstrate that ASE can estimate speeds reliably up to a distance of 3.5 m, with an average speed error of 0.08 m/s. When the object is close, the mean speed error is 0.016 m/s, which increases to 0.14m/s at 3.5m. Our system achieved an average detection rate of 99.0% in the case of direct walking, with a slight degradation to 95.0% at a distance of 3.5m.
@@ 5.3 Comparative Study
We compare ASE with three baselines: 1) SwiftTrack [89]: An acoustic speed algorithm for practical fast motion tracking, 2) DFS: A widely used method, implemented based on CAT [35], and 3) VeCare [88]: A recent work of statisitical acoustic sensing framework for child presence detection. The former two are DFS-based speed estimation approaches. For a fair comparison, we transmit modulated Zadoff-Chu (ZC) signals as in [89] with the same frame length, i.e., 10ms. We use unmixed CSI and extracted the frequency shift peak in the DFS spectrum for Doppler. We compare baselines under scenarios in §5.2, and also test the Non-Line-of-Sight (NLoS) condition. We apply the same settings as VeCare.
@@
Different Walk Means: Fig. 16, Fig. 16, and Fig. 16 compare the performance for direct walk, circle walk, and random walk, respectively. While ASE achieves a median error of 0.08m/s in direct walking, SwiftTrack and Doppler yield errors of 0.26m/s and 0.30m/s, respectively. Moreover, ASE significantly surpasses SwiftTrack and Doppler by 68.9% and 101% for circle walk and by 176.7% and 146.2% for random walk. We further depict the CDF of ASE against Doppler in Fig. 16. It shows DFS exhibits poor performance in circle walk while ASE retains consistent performance across various walking means. It is plausible for worse performance of SwiftTrack against Doppler in random walk, given its speed estimation algorithm is tailored for gesture tracking.
Figure 17: Impact of different directing angles
Figure 18: Impact of sequence lengths
Figure 19: Impact of different CSI rates
@@
Sensing Distance: A unique advantage of ASE is the enlarged sensing coverage, and we verify this by evaluation over different distances, as illustrated in Fig. 16, Fig. 16 and Fig. 16. While the error generally increases with respect to distances for all the methods, ASE consistently outperforms the baseline methods at all distances, with more significant performance gains at larger distances. As shown in Fig. 16, ASE achieves mean speed error less than 0.15m/s at 3.5 m, while SwiftTrack and Doppler soar to 0.42m/s and 0.43m/s, respectively. Overall, both SwiftTrack and Doppler degrade significantly beyond the distance of 1.5m. ASE obtains superior performance because our model leverages all multipath signals, which gets more observations across the room and expands the range to accommodate room-scale sensing.
@@
LoS vs. NLoS: We assess our system in NLoS settings with obstructions like a metal panel or a paper bag in front of the speaker and microphone. Despite sound’s known vulnerability to absorption [57], ASE remains effective in NLoS scenarios, exhibiting an average error rate of 0.13m/s and a standard deviation of merely 0.003m/s, as illustrated in Fig. 16. In contrast, the performance of SwiftTrack drops by 14.4% and 11.2% under metal and paper obstructions, respectively. We benefit from the effective use of multipath reflections, which is particularly helpful in NLoS scenarios.
@@
5.4 Benchmark Study
We present a micro-benchmark study of ASE in this section.
Figure 20: Impact of different interference sources
Figure 21: Impact of different sound volume level
Figure 22: Impact of amplitudes of sequence
Figure 23: Performance for different users
@@ 5.4 Benchmark Study
Sound Level: We modulate the transmission signal to the inaudible band to minimize interference with human hearing for daily use as detailed in Sec. 5.1, rendering the sound nearly inaudible. To evaluate the impact of different sound levels, we transmit signals at sound pressure levels of 33.4dB, 35.3dB, 36.7dB, 38.2dB, and 41.7dB, respectively. As shown in Fig. 23, the results indicate that the mean speed error decreases as the sound volume increases, as does the detection rate. Notably, our system’s maximum sound pressure of 41.7dB is significantly lower than most previous acoustic sensing studies [85, 88, 61, 37]. Despite the sound leakage problem on commodity devices [28], the sound noise of our system can be fairly neglected in commonly used settings.
Amplitude of Kasami Sequence: Besides volume of the device, the sound power is also determined by the amplitude of the original sequence. We hereby experiment with different amplitudes, shown in Fig. 23. This is not the final amplitude of the transmitted signal, as it must go through band modulation and PCM before transmission. As seen, the performance decreases first and then increases as amplitudes increase, and an amplitude of 2.5 strikes the best for both speed accuracy and detection rate. Amplitudes smaller than 2.5 will result in low sound power and thus large errors, while overlarge amplitudes will lead to distortions of the signal [28] due to the hardware limitation. In our case, sounds with amplitudes larger than 2.5 would be clipped off. Hence, we choose 2.5 as the amplitude of the original PN sequence.
@@
CSI Rate: We now study how CSI rates will impact the performance for different speeds. To control the moving speed, we use a programmable rail track with a plate on it. We test with speeds of 0.3m/s to 1.6m/s, and compare the performance with a CSI rate of 94 Hz and 188 Hz, respectively. We apply OTDM scheme to achieve the CSI rate of 188 Hz. As shown in Fig. 19, a low CSI rate cannot measure large speeds and will produce large errors. The mean errors for speeds of 0.6m/s and 0.8m/s with a CSI rate of 94Hz are 0.18m/s and 0.44m/s, respectively. Comparatively, the corresponding errors with that of 188Hz are 0.055m/s and 0.066m/s, respectively. Notably, our system maintains a relatively low average error of 0.11m/s even at 1.6m/s. Overall, the results clearly demonstrate the need for a sufficient CSI rate for speed estimation and justify the effectiveness of the proposed OTDM scheme. In ASE, we can hold speed up to 1.6m/s using single-speaker, adequate for capturing indoor walk speed.
@@
Different Subcarrier Combining Weight: We test different subcarrier combing weight method. Our results show using prominence as the weight leads to a mean speed error of 0.13m/s, whereas MS results in an error of 0.19m/s. Besides, we compare different weight decaying algorithms. We observe a 15% decrease in mean speed error when applying sigmoid weight decay. These results confirm the effectiveness of our subcarrier combination weight in ASE.
Different Location: To test ASE in different environments, we perform experiment in different locations, including a conference room and an office. The experiment subject is walking randomly within 4m range. We have achieved 100% and 97.13% detection rate in these scenarios. Therefore, our system can adapt to different indoor environments.
Different Users: We also examine how ASE performs for different users. We recruit four volunteers and let them walk freely with varying speeds, orientations, and trajectories. The results are shown in Fig. 23, which shows a mean error of 0.12m/s and an average detection rate of 97.4%, with only marginal differences among different users.
Figure 24: Speed profile of different human activities
Figure 25: Speed profile of continuous activities
Figure 26: Gait step and Cycle Estimation
System Overhead: We mainly employ MacBook 2021 to run the experiments. Our statistical algorithms are lightweight, with an average latency of 2.32s over 40s acoustic data. The channel estimation takes 0.20s on average and the weight decay and combination accounts for 0.41s.
5.5 Case Study
We present three case studies to showcase various ASE-enabled applications.

### v2

@@ 4.2. Transmission Scheme
In this part, we will elaborate on our transmission scheme design. We will then describe the channel estimation principles.
@@
Single-Speaker OTDM: We stack two orthogonal signals as a complex signal and pass it into the I/Q modulator for transmission. Specifically, we can treat the orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$ as the real and imaginary parts of a complex signal, i.e., $s(t)=s_{1}(t)+js_{2}(t)$. To delay $s_{2}(t)$, we shift the sequence by padding zeros to the front of $s_{2}(t)$. Practically, however, a speaker can only transmit a single sequence of real signals. To transmit the resultant complex signal over a speaker, we pass it through an I/Q modulator to convert it into a real one. As illustrated in Fig. 6, we additionally multiply $s_{1}(t)$ and $s_{2}(t)$ with a carrier signal of orthogonal phase and obtain the modulated real signal: $s(t)=s_{1}(t)\cos(2\pi f_{c}t)-s_{2}(t)\sin(2\pi f_{c}t)$. With such modulation, the two components remain orthogonal to each other, while the modulated signal can be transmitted through a single speaker. At Rx, the received signal is first demodulated by the carrier signal with a low-pass filter (LPF). We can then separate the overlapped signal and estimate the CSI for each sequence at the receiver’s side, as shown in Fig. 6. By doing so, we can obtain two separated signals $r_{1}(t)$ and $r_{2}(t)$. Then, channel estimation will be performed as follows.
@@
Figure 8. Peak prominence weighting examples. Top: Violin plots of the distribution of the delays of the first peak. Middle: Original weight distribution, where the circle size indicates the weights. The red dashed line represents the fitted $\sigma(\tau)$. Bottom: Adjusted prominence weights.
@@
Channel Estimation: To capture the channel properties, we generate the Kasami sequence $s(t)$ of length $N_{s}$ as the probe signal and transmit it over a speaker. The signal interacts with the environment before arriving at the microphone, where it undergoes scattering in the sound diffusion field. At the receiver end, we acquire $r(t)$ and separate it into $r_{1}(t)$ and $r_{2}(t)$, respectively. The CIR $h(t)$ is then estimated as the correlation between them, i.e., $h_{i}(t)=r_{i}(t)\ast s(t),i\in\{1,2\}$. Then CSI $H_{i}(f,t)$ is obtained by transforming $h_{i}(t)$ into the frequency domain via FFT. $H_{1}(f,t)$ and $H_{2}(f,t)$ will be then merged using the OTDM scheme to acquire the boosted channel estimation $H(f,t)$, as illustrated in §2.
Synchronization between Microphone and Speaker Pairs: While the microphone and speaker are connected to the same controller and should theoretically be temporally aligned, hardware imperfections can introduce small time discrepancies between the Tx and Rx pair. Traditional FMCW tracking algorithms usually require precise synchronization to correctly determine the start time of IF signals. However, our approach minimizes the impact of synchronization issues. At a high level, any discrepancies manifest as phase offsets in the asynchronized Channel State Information (CSI). This is not a significant problem for ASE because we do not introduce blank intervals and instead focus on extracting channel power to determine speed. Specifically, the asynchronous CSI would be
(15) $H_{d}(f)=H(f)\cdot\exp(-j\cdot 2\pi f\cdot\tau)$
where $\tau$ is the temporal offset. And the channel power, as denoted in (8) can be written as
(16) $G_{d}(f)$ $=|H_{d}(f)|^{2}+\epsilon(f,t)$
$=|H(f)|^{2}+\epsilon(f,t)=G(f),$
which is equivalent to the synchronized version. Therefore, the temporal offsets do not affect the magnitude of the channel power, ensuring reliable speed acquisition even without precise synchronization.
4.3. Speed Estimation
@@
(a) Overall Setting.
(b) Different Environments.
Figure 10. Experiment Setting.
Weighted Subcarrier Combining: Proposed model accounts for all multipath components for speed estimation. Recall Eq. (13), however, the theoretical model estimates speed from a single subcarrier. Given multiple subcarriers available, we can further embrace frequency diversity to boost speed estimation. Instead of performing speed estimation on each subcarrier and fusing the estimates, we propose to properly combine $\psi_{G}(f,t)$ calculated on each subcarrier to enhance the SNR of the speed signal. Formally, we aim to obtain the boosted ACF as $\hat{\psi}_{G}(\tau)=\sum w(f)\psi_{G}(f,\tau)$, where
[…]
@@
Frequency Alignment: The weighted combining will enhance the speed signals and reduce the noise, only if the signals are synchronized. However, significant subcarrier frequency variations misalign the ACFs across subcarriers, particularly the first peak, for the same speed. Take two subcarriers on 20 kHz and 24 kHz as an example. Considering a speed $v=0.5m/s$, the peaks will appear at a delay $\tau=x_{0}\lambda(f)/2\pi v$. The wavelength $\lambda(f)$ depends on the carrier frequency, and will be 1.715 cm for 20 kHz and 1.429 cm for 24 kHz, leading to two considerably different delays for the first peaks in the respective ACFs. Therefore, we need to eliminate the subcarrier frequency differences and align all the ACFs. We scale the ACF in the lag domain and align them with respect to the first peaks. Particularly, we choose a virtual frequency as a reference, denoted as $f_{ref}$, and scale the ACF on all subcarriers to the same $f_{ref}$, i.e., $\psi(f,\tau)\to\psi(f_{ref},\tau^{\prime})$. Recall Eq. (14), we obtain the aligned first peak as,
(18) $v(f,\tau_{s})=\frac{x_{0}c}{2\pi f\tau_{s}}=\frac{x_{0}c}{2\pi f_{ref}\tau_{s}^{\prime}}=v(\tau_{s}^{\prime}),$
where $\tau_{s}^{\prime}=\frac{f}{f_{ref}}\tau_{s}$ represents the aligned first peak. Practically, it is implemented by interpolation. Then the weighted combining can be done on the aligned ACF.
Figure 11. MSE-DW.
Figure 12. MSE-CW.
Figure 13. MSE-RW.
5. Evaluations
5.1. Implementation
@@
Hardware: We implement ASE with programmable smart speaker prototype, i.e., a UMA-8-SP USB microphone (USB, 2023), and AS05308AS-R speaker(Spe, 2022). Notably, we only use one microphone channel in our experiments. The microphone and the speaker are co-located. The hardware is connected to a power hub and MacBook Pro 2021.
@@
Software: We implement the algorithms of ASE in Python and Matlab. Specifically, we develop a sound playing and recording program with Python and implement the pipeline of processing algorithms with MATLAB. We apply a 1-s window to compute ACF with a step of 0.1s. We perform motion detection, based on ZCC, to determine whether a user is moving and perform speed estimation when motion is detected. Before weighted subcarrier combining, we also perform outlier detection to sift out some abnormal ACFs, which occasionally appear as either a near-linear trend or a zig-zag spike.
@@
Data Collection: As shown in Fig. 10, we mainly conduct our experiment in a 4 $\times$ 4m room equipped with the OptiTrack (Mot, 2023), which serves as the ground truth. It is a camera-based precise motion capture system. We derive speed from the solved whole-body skeleton coordinates from OptiTrack. To calibrate and fuse the data across cameras, participants wear specialized clothes with visual markers. The experiment is conducted in a regular office environment and is subject to various background noises, including footsteps in the corridor and the sound of the room air conditioner. During the experiment, only the test participant is walking while others in the room are stationary. We have gained IRB approval from our university for data collection. In total, we have collected over 700 minutes of moving data traces. We evaluate the performance of ASE under various settings.
Metrics: We use Mean Speed Error (MSE) between the estimated speed and the ground truth and detection rate (i.e., how often ASE can reliably detect a speed) as the main evaluation metrics. Our results demonstrate the effectiveness of our approach in accurately estimating speed and detecting movement in a daily living environment. We present detailed results below.
Figure 14. Detection Rate.
Figure 15. CDF vs. Distances.
Figure 16. CDF vs. Walk Means.
Figure 17. MSE of NLOS Case.
5.2. Overall Performance
@@
Direct Walk (DW): We place the device at the center of one wall in the room. Participants are asked to walk towards the audio device from 1m to 3.5m. We then calculate the mean speed error in Fig. 13 and detection rate in Fig. 17. Our results demonstrate that ASE can estimate speeds reliably up to a distance of 3.5 m, with an average speed error of 0.08 m/s. When the object is close, the mean speed error is 0.016 m/s, which increases to 0.14m/s at 3.5m. Our system achieved an average detection rate of 99.0% in the case of direct walking, with a slight degradation to 95.0% at a distance of 3.5m.
@@ 5.3. Comparative Study
We focus on acoustic-based speed estimation and compare ASE with three baselines: 1) SwiftTrack (Zhang et al., 2023b): An acoustic speed algorithm for practical fast motion tracking, 2) DFS: A widely used method, implemented based on CAT (Mao et al., 2016), and 3) VeCare (Zhang et al., 2023a): A recent work of a statistical acoustic sensing framework for child presence detection. The former two are DFS-based speed estimation approaches. For a fair comparison, we transmit modulated Zadoff-Chu (ZC) signals as in (Zhang et al., 2023b) with the same frame length, i.e., 10ms. We use unmixed CSI and extracted the frequency shift peak in the DFS spectrum for Doppler. We compare baselines under scenarios in §5.2, and also test the Non-Line-of-Sight (NLoS) condition. We apply the same settings as VeCare. We will first compare SwiftTrack and Doppler and leave VeCare as the last part.
@@
Different Walk Means: We compare ASE with SwiftTrack and Doppler in three means: direct walk, circle walk and random walk. Our results are shown in Fig. 13, Fig. 13 and Fig. 13 respectively. While ASE achieves a median error of 0.08m/s in direct walking, SwiftTrack and Doppler yield errors of 0.26m/s and 0.30m/s, respectively. Moreover, ASE significantly surpasses SwiftTrack and Doppler by 68.9% and 101% for circle walk and by 176.7% and 146.2% for random walk. We further depict the CDF of ASE against Doppler in Fig. 17. It shows DFS exhibits poor performance in circle walk while ASE retains consistent performance across various walking means. It is plausible for worse performance of SwiftTrack against Doppler in random walk, given its speed estimation algorithm is tailored for gesture tracking.
Figure 18. Impact of different directing angles.
Figure 19. Impact of sequence lengths.
Figure 20. Impact of different CSI rates.
@@
Sensing Distance: A unique advantage of ASE is the enlarged sensing coverage, and we verify this by evaluation over different distances, as illustrated in Fig. 13, Fig. 13 and Fig. 13. While the error generally increases with respect to distances for all the methods, ASE consistently outperforms the baseline methods at all distances, with more significant performance gains at larger distances. As shown in Fig. 13, ASE achieves a mean speed error less than 0.15m/s at 3.5 m, while SwiftTrack and Doppler soar to 0.42m/s and 0.43m/s, respectively. Overall, both SwiftTrack and Doppler degrade significantly beyond the distance of 1.5m. ASE obtains superior performance because our model leverages all multipath signals, which get more observations across the room and expand the range to accommodate room-scale sensing.
@@
LoS vs. NLoS: We assess our system in NLoS settings with obstructions like a metal panel or a paper bag in front of the speaker and microphone. Despite sound’s known vulnerability to absorption (Tang and Yan, 2017), ASE remains effective in NLoS scenarios, exhibiting an average error rate of 0.13m/s and a standard deviation of merely 0.003m/s, as illustrated in Fig. 17. In contrast, the performance of SwiftTrack drops by 14.4% and 11.2% under metal and paper obstructions, respectively. We benefit from the effective use of multipath reflections, which is particularly helpful in NLoS scenarios.
@@
5.4. Benchmark Study
In this section, we present a detailed analysis of various impact factors in ASE. We begin by investigating the impact of different movement directions on the accuracy of our approach. Next, we examine the performance of our system extending to dual speakers to conduct the experiment. We also study the impact of various acoustic factors, including sound level, sound amplitude, and interference sources.
Figure 21. Impact of different interference sources.
Figure 22. Impact of different sound volume levels.
Figure 23. Impact of amplitudes of original sequence.
@@ 5.4. Benchmark Study
Sound Level: We modulate the transmission signal to the inaudible band to minimize interference with human hearing for daily use as detailed in Sec. 5.1, rendering the sound nearly inaudible. To evaluate the impact of different sound levels, we transmit signals at sound pressure levels of 33.4dB, 35.3dB, 36.7dB, 38.2dB, and 41.7dB, respectively. As shown in Fig. 23, the results indicate that the mean speed error decreases as the sound volume increases, as does the detection rate. Notably, our system’s maximum sound pressure of 41.7dB is significantly lower than most previous acoustic sensing studies (Yun et al., 2017; Zhang et al., 2023a; Wang et al., 2019b; Mao et al., 2018). Despite the sound leakage problem on commodity devices (Li et al., 2022a), the sound noise of our system can be fairly neglected in commonly used settings.
Amplitude of Kasami Sequence: Besides the volume of the device, the sound power is also determined by the amplitude of the original sequence. We hereby experiment with different amplitudes, shown in Fig. 23. This is not the final amplitude of the transmitted signal, as it must go through band modulation and PCM before transmission. As seen, the performance decreases first and then increases as amplitudes increase, and an amplitude of 2.5 strikes the best for both speed accuracy and detection rate. Amplitudes smaller than 2.5 will result in low sound power and thus large errors, while overlarge amplitudes will lead to distortions of the signal (Li et al., 2022a) due to the hardware limitation. In our case, sounds with amplitudes larger than 2.5 would be clipped off. Hence, we choose 2.5 as the amplitude of the original PN sequence.
@@
CSI Rate: We now study how CSI rates will impact the performance for different speeds. To control the moving speed, we use a programmable rail track with a plate on it. We test with speeds of 0.3m/s to 1.6m/s, and compare the performance with a CSI rate of 94 Hz and 188 Hz, respectively. We apply the OTDM scheme to achieve the CSI rate of 188 Hz. As shown in Fig. 20, a low CSI rate cannot measure large speeds and will produce large errors. The mean errors for speeds of 0.6m/s and 0.8m/s with a CSI rate of 94Hz are 0.18m/s and 0.44m/s, respectively. Comparatively, the corresponding errors with that of 188Hz are 0.055m/s and 0.066m/s, respectively. Notably, our system maintains a relatively low average error of 0.11m/s even at 1.6m/s. Overall, the results clearly demonstrate the need for a sufficient CSI rate for speed estimation and justify the effectiveness of the proposed OTDM scheme. In ASE, we can hold speed up to 1.6m/s using single-speaker, adequate for capturing indoor walk speed.
Figure 24. Different Height.
Figure 25. Different Users.
Figure 26. Different Environment and Locations.
@@
Different Subcarrier Combining Weight: We test different subcarrier combining weight methods. Our results show that using prominence as the weight leads to a mean speed error of 0.13m/s, whereas using motion statistics simply (Zhang et al., 2023a) results in an error of 0.19m/s. Besides, we compare different weight decaying algorithms. We observe a 15% decrease in mean speed error when applying sigmoid weight decay. These results confirm the effectiveness of our subcarrier combination weight in ASE.
Different Device Height: To investigate the impact of varying the height of the transmitter and receiver pair on system performance, we conducted experiments using ASE with different hardware heights, as shown in Figure 26. By utilizing a tripod, we controlled the height range from 80 to 115 cm. The results indicate that ASE retains consistent accuracy under changes in heights, with a mean of 0.13 m/s and a standard deviation of 0.002 m/s. Despite these variations, the average detection rate remains high at 99.8%. This suggests that the height of the devices does not significantly affect the overall performance of ASE.
Different Users: We also examine how ASE performs for different users. We recruit four volunteers and let them walk freely with varying speeds, orientations, and trajectories. The results are shown in Fig. 26. Our approach achieves 80% tile error of 0.2m/s. The detection rates of three users are 95.55%, 98.95% 95.03% and 100%, respectively. Accordingly, the mean error of four users is 0.12m/s with an average detection rate of 97.4%. These findings demonstrate the effectiveness and accuracy of ASE in estimating the speed of human movement in indoor environments. Our approach achieves high detection rates and low error rates across multiple users with varying movement patterns and speeds. These results have important implications for the use of ASE in real-world applications, such as surveillance, security, and healthcare, where accurate estimation of human movement speed is critical.
Different Environments and Locations: To evaluate ASE’s robustness across various settings, we conduct experiments in multiple environments, as illustrated in Fig. 10(b). We deploy devices in a lab, office, conference room, hallways, kitchens, and an open space. These locations differ in geometry, building materials, floor heights, and ambient noise levels. Except for the lab and office, we leverage the depth camera to get the range information and then convert it to speed. During the tests, participants walk randomly within each environment. We compute the mean speed error and detection rate for all scenarios, as summarized in Fig. 26. In indoor environments, the average speed error is 0.136 m/s with a standard deviation of 0.01 m/s, which aligns with previous benchmark results. The corresponding detection rate averages 0.998, with a standard deviation of 0.003. These results demonstrate that ASE consistently performs well across varied indoor settings, proving its robustness. We also evaluate ASE in an open space, where the mean speed error increases to 0.21 m/s, although the detection rate remains high. This outcome is expected given the relative scarcity of reflectors in open environments, which can impact speed estimation accuracy. Nevertheless, the consistently high detection rate highlights the robustness of our motion detection algorithm. Future work can explore the development of advanced speed estimation methods tailored for open areas.
Figure 27. Speed profile of different human activities.
Figure 28. Speed profile of continuous activities.
Figure 29. Gait step and Cycle.
Runtime Analysis: In this section, we evaluate the runtime performance on various platforms. We first test the algorithm’s performance on a MacBook 2021 (with Apple M1 Pro chip), analyzing a 10-second audio segment using a 1-second sliding window with a 0.1-second step. The algorithm demonstrates exceptional efficiency, achieving a total runtime of 2.299 seconds, well below the 10-second data duration, which confirms its capability for real-time processing. The decoding of OTDM takes 0.0999 seconds, and the ACF part finishes in 0.831 seconds. Peak prominence and weight computations are notably fast at 0.098 seconds and 0.067 seconds, respectively. Additionally, memory usage is a modest 387.56 MB, well within the capacity of modern systems, making this algorithm highly suitable for efficient, real-time processing. Additionally, we deploy our system on an embedded device, i.e., Raspberry Pi Compute Module 4, which is equipped with 4-core ARM-v8 processors. For the same 10-second audio segment with identical windowing parameters, the algorithm processes the respiration data with an overall runtime of 3.912 seconds, still significantly faster than the data duration, ensuring real-time performance even on resource-constrained hardware. The decoding takes 0.384 seconds, and ACF calculations total 0.440 seconds. Memory usage on the Raspberry Pi is optimized at 318.57 MB, demonstrating the algorithm’s ability to operate efficiently within the limited resources of an embedded system. These results highlight the algorithm’s robustness and adaptability, enabling real-time speed analysis across both high-performance and embedded platforms.
5.5. Case Study
In this section, we present a case study on three applications of human walking speed: activity recognition, fall detection, and gait recognition, to showcase various ASE-enabled applications.